We performed a set of tappings in order to asses the status of scattered light noise at the WE after the installation of the new WE mirror with the front coated baffle in place. We tapped the same positions done in 65378. Specifically:
- positions on CryoTrap side: CT holding points (Nort-East, North-West, South-East) , VAT flange
- positions HR side: bellow, 1200 flange, cylindric vacuum pipe
- positions AR side: cylindric vacuum pipe, 1200 flange, viewport
All times of tappings and notes are listed in the attached file. Each tappping lasted 120seconds.
Some observations we made:
The CryoTrap side is much quiter than before, in fact despite providing approx. the same level of seismic excitation, we do not see significant impact in Hrec (just a tiny peak at 195Hz when tapping on the SE holding point) and associated to a peak in CT accelometer. See Figure 1. Also, tiny peaks around 100Hz and at 195Hz when tapping on the VAT flange. See Figure 2.
The position on the tower at the HR side cause some some excitation in hrec around 100Hz, a bit louder when tapping on the 1200 flange. When tapping close to the bellow, in addition to small noise structures around 100Hz we also caused the rise of hrec noise below 30Hz: Figure 3 and Figure 4 (spectrogram showing the set of the three tappings HR side)
The AR side instead looks more touchy. Tappings on AR side excited similar peaks as Before (66031). In Particular at 90Hz, 167 Hz and 42 Hz (narrow). We observe that the peak at 167Hz looks a resonance of the flange, since it gets very loud in the accelerometer (AR flange) in which a residual is present also during the quiet, see Figure 5. When tapping on the viewport we notice a similar correspondence between the 90Hz peak in Hrec and the accelerometer: Figure 6.
The overall impression is that while we mitigated the sensitivity at the CT side, the AR side remains roughly the same. Some scattering (low level) seems to happen at the AR side, maybe the flange.
Figure 7 is an overview.